OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given = 5 sqrt 6

OpenStudy (anonymous):

thnx:)

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

Let the base be x. The height will also be x since this is an isosceles triangle. This is also a right triangle, so we can apply Pythagoras Theorem: x^2 + x^2 = (5sqrt(6))^2 2x^2 = (25)(6) x^2 = (25)(3) x = sqrt((25)(3)) x = 5sqrt(3) Hence, leg of triangle is 5sqrt(3).

OpenStudy (anonymous):

Lol I accidentally deleted the problem. Oops

OpenStudy (anonymous):

thnx lol

OpenStudy (anonymous):

u helped a lot:))